# LD-SDM: Language-Driven Hierarchical Species Distribution Modeling

### v1

@@ Abstract
We focus on the problem of species distribution modeling using global-scale presence-only data. Most previous studies have mapped the range of a given species using geographical and environmental features alone. To capture a stronger implicit relationship between species, we encode the taxonomic hierarchy of species using a large language model. This enables range mapping for any taxonomic rank and unseen species without additional supervision. Further, we propose a novel proximity-aware evaluation metric that enables evaluating species distribution models using any pixel-level representation of ground-truth species range map. The proposed metric penalizes the predictions of a model based on its proximity to the ground truth. We describe the effectiveness of our model by systematically evaluating on the task of species range prediction, zero-shot prediction and geo-feature regression against the state-of-the-art. Results show our model outperforms the strong baselines when trained with a variety of multi-label learning losses.
@@ 1 Introduction
The ever-increasing availability of crowdsourced and remote sensing data has opened opportunities to solving challenging tasks that were not possible before. One such task is global-scale joint species distribution modeling (SDM). Understanding species ranges is not only essential for biodiversity conservation but also for understanding the effects of climate change. However, complex biophysical and geographic processes make SDM a challenging task.
Traditionally, species ranges were mapped using geographical and environmental features. While the current approaches have their merits, we contend they fall short in accurately representing the inter-species relationships inherent in the taxonomic hierarchy. For example, species belonging to the same genera tend to be found in similar locations. They often share many common characteristics, including their general habitat preferences [41]. Nevertheless, this pattern diminishes as one moves up in the taxonomic hierarchy. To capture these complex hierarchical relationships between the species, we use large language models (LLMs) to encode their taxonomic information in text (as shown in Figure 1). Intuitively, this should enhance the identification of rare taxa [9]. To the best of our knowledge, no previous work has incorporated large language models for SDM.
Figure 1: Species Distribution Modeling using LLM Guidance. We use text prompts in the form of key-value pairs for training our species distribution model. This kind of prompt captures the taxonomy hierarchy of species. We use only the exhaustive prompt (d) for training while the rest are used for zero-shot predictions.
In order to incorporate species-specific embeddings into SDM, it is necessary to reformulate the SDM problem. Previous approaches to SDM [27, 15, 24] tried to answer the question, “which species are likely to be found at a given location?”. In contrast, our approach to SDM models, “which locations are likely for a given species to be found?”. Assuming uniform prior over the species distribution [27, 24], this is an equivalent formulation as before. But by doing so, we can now incorporate a variety of information, including species-specific embeddings or any other metadata, which opens up new opportunities for more accurate modeling. Alternatively, this formulation makes it easier to incorporate unseen species that were not present in the training data.
@@
When it comes to predicting species range, the evaluation process typically involves using range maps that are defined by experts or curated carefully [29, 43]. However, these maps can often be sparse and difficult to obtain. Alternatively, using any crowdsourced species observation data can also be challenging as there is no precise definition of species absence. Intuitively, models with false predictions close to their true locations are better than those with distant false predictions. This is because models in the former case learn spatially localized species distributions. Previous studies have used the area under the curve (AUC) [29, 43] and the mean average precision (MAP) [15] to evaluate SDMs. However, these metrics ignore the spatial distribution of predictions and weigh them equally, making them highly sensitive to spatial shifts in the predictions. Even a single pixel-level shift in predictions could add significant bias to the scores.
To this end, we propose a proximity-adjusted evaluation metric that penalizes predictions based on their distance to the closest true locations. We construct a One-Way Probability-Weighted Chamfer Distance (PWCD) metric, that measures the dissimilarity between the true and predicted species range maps, which is robust to small shifts. This enables an effective evaluation of SDMs using pixel-level range maps and crowdsourced data (where absence is not confirmed).
The contribution of our work is threefold:
@@ 2 Related Works
Species distribution modeling (SDM) is a set of techniques that uses environmental, geographic, and species observation data to predict species distribution across geographic space and time [16, 4]. In this context, we restrict ourselves to the task of species range mapping over geographic space only. In the past, SDM relied on hand-crafted geographic and environmental features as input to capture the implicit spatial correlations in data [17, 29, 43]. These approaches were limited to one species per model due to their reliance on traditional supervised learning algorithms. With advancements in deep neural network architectures, end-to-end frameworks are now able to model joint species distributions and complex implicit relations in data [6, 10, 15].
Another important aspect of SDM is the type of species occurrence data. Species occurrence data may exist in one of the two forms: presence-only and presence-absence. Collecting presence-absence observations is a challenging task, requiring expert surveys especially to confirm the absence of a given species in some region [28]. However, once collected, this kind of data can be easily fed into supervised learning frameworks, without requiring sophisticated modifications [19, 11, 20, 6, 40]. With no expert knowledge, presence-only data can be recorded when any species of interest is encountered. This way of collecting data does not require the data collector to confirm absences. However, since there are no negative samples, it is difficult to use standard supervised learning frameworks [10, 22, 8].
Most of the recent works formulate the problem of SDM using presence-only data as a multi-label classification task [27, 15]. Here, the goal of SDM is to predict the likelihood over species for a given location and environmental features. A variety of multi-label learning losses have been proposed in the literature that can easily be plugged in for this task. For instance, the AN-full (full assume negative) loss [27] randomly samples locations that are assumed to be negative during training. The ME-full (full maximum entropy) [49] tries to maximize the entropy of predictions of negative samples and at randomly sampled locations during training. Other losses such as ASL (asymmetric loss) [32], Hill loss [48], and RAL (robust asymmetric loss) [31] weigh the negative samples in a way that easy negative and potential false negative samples are down-weighted. The authors of [15] presented an extensive comparison of the performance of SDMs trained with losses involving pseudo-negative sampling. Here, we extend their work by additionally comparing their SDMs trained with the ASL and RAL losses.
@@
Finally, another line of work uses high-resolution satellite imagery at each species occurrence location to guide SDM methods [7, 13, 25, 26, 35, 39]. For example, [39] uses satellite imagery and geolocation to directly predict the likelihood over species while [35] uses a contrastive learning framework to establish similarity between ground-level images of species and corresponding satellite imagery. These methods are limited in geographic scale due to the high-resolution nature of raw data [15]. Except [39], all the works view this problem as an information retrieval task rather than a species prediction task.
@@ 3.1 Problem Definition
We consider presence-only data in the form of occurrence records: $\mathcal{D}=\{(l_{i},y_{i})|\>i=1,...,N\}$, where $l_{i}\in\mathbb{R}^{2}$ is the geolocation of each observation and $y_{i}\in[1,...,S]$ is the species of the observed organism. Additionally, a set of environmental features $e_{i}\in\mathbb{R}^{n}$ is included for each location $l_{i}$. Many previous studies [27, 24, 15] have tackled the task of SDM by training a neural network to model the conditional probability of $p(y|\>l,e)$. This translates to the problem of “how likely a species can be found at a given location”. However, our approach involves modeling the probability of $p(r|\>y,l_{r},e_{r})$, which answers the question “Is it likely for a given species to be found at given location?”. Here, $r\in[1,...,L]$ is a set of discrete locations that are of interest when predicting the range of species belonging to some class. In this framework, we define the set of locations $r$ as a contiguous grid that is laid over the surface of the globe. The variables $l_{r}$ and $e_{r}$ represent fixed parameters in the model, namely the location vectors and environmental features defined at each location in $r$ respectively. It is easy to see that the input $y$ and the parameters $(l_{r},e_{r})$ can have any arbitrary representations. So, this new formulation allows for flexible encoding of both the species class and geographic location, compared to the previous formulation which only allowed flexibility in representing the geographic location.
@@ 3.2 Architecture Overview
Our objective is to train a neural network of the form: $f_{\theta}(y;l_{r},e_{r}):\mathbb{R}^{d}\rightarrow[0,1]^{L}$, where $d$ is the dimensionality of the species feature vector. $f_{\theta}$ is composed of three modules: 1) a location encoder network $f_{\alpha}$, 2) a species encoder network $g_{\beta}$, and 3) a multi-label classifier $h_{\gamma}$. See Figure 2 for a high-level architecture diagram. The network $f_{\alpha}$ consists of convolutional feature extractors and the Spherical Fourier Neural Operator (SFNO) [5] blocks that take as input the fixed parameters: $l_{r}$ and $e_{r}$. The location vectors are used as learnable positional encodings for the SFNO blocks. The species encoder $g_{\beta}$ takes as input the text descriptions corresponding to a particular species of class $y$. The multi-label classifier $h_{\gamma}$ consists of a cross-attention network that fuses features from the SFNO and text encoder and performs multi-label classification using a linear layer. We call our model Language-Driven Species Distribution Model (LD-SDM).
@@ 3.3 Species-specific embeddings
Many popular crowdsourcing science platforms, such as iNaturalist [44], contain not just occurrence records, but also valuable species-specific information. This information may include the complete taxonomic hierarchy of each species (such as class, order, family, etc.) or the IUCN red list category (such as endangered, near threatened, etc.). While often ignored, this fine-grained information could be incorporated to enhance the performance and interpretability of SDMs. In this work, we incorporate the taxonomic hierarchy of species into our SDM in the form of text. Each species is uniquely identified by its taxonomic classification ranks, which are represented as key-value pairs. For instance, a key-value pair that represents the wild duck is shown below:
t = {class: Aves, order: Anseriformes, family: Anatidae, genus: Anas, species: Anas platyrhynchos}
Recently, Large Language Models (LLMs) have demonstrated excellent capabilities in representing implicit knowledge, linguistic structure, and hierarchical concepts when trained on massive amounts of data [1]. With no finetuning, pretrained LLMs have shown impressive zero-shot performance with zero in-domain data. In this work, we use the recently released, LLaMA-2 [42] to encode the species-specific key-value pairs as text. We experiment with three variants of LLaMA-2: 7B, 13B, and 70B.
@@ 3.4 Loss Functions
Our formulation of the SDM falls in the domain of multi-label long-tailed learning, where the model learns to predict the likelihood probability over each location for a given species. Since we are modeling the likelihood over locations, our problem differs from the Single Positive Multi Label (SPML) task [14]. Here, we are given multiple positive locations of occurrence for each species. We explore different types of losses for our models that do not require pseudo-negative sampling, such as ASL and RAL. However, for the state-of-the-art models, we use all types of losses, including AN-full and ME-full, thereby extending their work. We describe the losses in detail below.
@@
$\mathcal{L}_{\text{AN-full}}(\hat{y},y)=-\frac{1}{S}\sum_{s=1}^{S}[\mathbbm{1}_{[y_{s}=1]}\alpha log(\hat{y}_{s})$ (1)
@@
$\mathcal{L}_{\text{ME-full}}(\hat{y},y)=-\frac{1}{S}\sum_{s=1}^{S}[\mathbbm{1}_{[y_{s}=1]}\alpha log(\hat{y}_{s})$ (2)
@@
where $\psi(\hat{y}_{s})=\lambda-\hat{y}_{s}$ is the Hill loss term. $M$, $N$, $\alpha_{n}$ and $\beta_{n}$ are balance parameters for positive and negative samples.
@@ 4.1 Implementation Details
As described in section 3.2, we use SFNO to extract geographic and environmental-specific features. The architecture is borrowed from [5]. We use two spectral and encoder layers in the SFNO. The encoder dimension used is 128 while the number of SFNO blocks are two. For the coordinate-only models, we drop the environmental features during the training. For the text encoder, we use frozen LLaMA-2 [42] (70B variant) and extract its encoder outputs for each species. Before training, we precompute and save the embeddings for each species in the disk. For the multi-label classifier, we use a single-layer multi-headed cross-attention network followed by a fully-connected layer with sigmoid activations. We compare LD-SDM with two state-of-art-models: SINR [15] and SIREN(SH) [34]. SINR learns an end-to-end neural representation of geographic locations that is used to predict the likelihood over species. SIREN(SH) uses a spherical harmonic representation of geographic locations which is then used to predict the likelihood over species. The rest of the details about our models and the implementation of state-of-the-art models are described in Appendix C.
@@ 4.2 Proximity-Aware Evaluation
Evaluating SDMs is a challenging task since it requires the comparison of spatially distributed probability values. Pixel-wise comparison methods are not well-suited for this task since they are susceptible to spatial prediction shifts. Intuitively, false positive predictions that are geographically distant from true positive predictions are more significant than those that are close. This is not inherent in the evaluation metrics used by previous studies such as the Kullback–Leibler divergence, AUC, or MAP, which are invariant of the spatial arrangement of probability mass.
@@
A well-known spatially explicit distance used for comparing any two distributions is the Earth-Mover (Wasserstein) Distance. Due to its computational requirements and infeasibility on large-scale data points, Chamfer Distance is a popularly used proxy [3]. However, it has only been developed for point clouds and there doesn’t exist a chamfer distance measure to compare two spatial probability distributions. Inspired by [46], we make a simple modification to their Density-Aware Chamfer Distance (DCD) metric that works for comparing a predicted probability distribution map ($P_{1}$) with a binary probability map ($P_{2}$) representing the ground truth. We add a likelihood probability term which weights the distance term. We call this metric One-Way Probability Weighted Chamfer Distance (PWCD), which measures the false positive rate in a proximity-aware manner. This metric is between [0,1] and could also be used to compute the number of false positives instead of the rate. It is given by:
$d_{\text{PWCD}}(P_{1},P_{2})=\frac{1}{|N|}\sum_{x\in P_{1}}\left(1-e^{-\alpha p(x)||x-y||_{2}}\right)$ (5)
$FP_{\text{PWCD}}(P_{1},P_{2})=\sum_{x\in P_{1}}\left(1-e^{-\alpha p(x)||x-y||_{2}}\right)$ (6)
where, $y=\underset{z\in P_{2}\cap p(z)=1}{argmin}||x-z||_{2}$. This represents the nearest neighbor ground-truth observation of a given pixel $x$ in the predicted probability distribution map. $N=\{z\in P_{2}|\>p(z)=0\}$ is the set of all negative observations in the ground truth. $\alpha$ is a temperature parameter that controls the exponential weighting of the distance. See Appendix B for more details. For our case, $|P_{1}|=|P_{2}|=L$. Notice that we have omitted the density term in the metric as our probability maps are a contiguous grid of pixels and equally dense everywhere. Since it is a proxy for false positives, it can be plugged into any other metric of choice such as the MAP and AUC. In this work, we report proximity-adjusted MAP scores denoted as MAP(pa) by replacing false positives with our metric. All the scores reported are an average over scores computed for each species category.

### v2

@@ Abstract
We focus on species distribution modeling using global-scale presence-only data, leveraging geographical and environmental features to map species ranges, as in previous studies. However, we innovate by integrating taxonomic classification into our approach. Specifically, we propose using a large language model to extract a latent representation of the taxonomic classification from a textual prompt. This allows us to map the range of any taxonomic rank, including unseen species, without additional supervision. We also present a new proximity-aware evaluation metric, suitable for evaluating species distribution models, which addresses critical shortcomings of traditional metrics. We evaluated our model for species range prediction, zero-shot prediction, and geo-feature regression and found that it outperforms several state-of-the-art models.
@@ 1 Introduction
Species distribution modeling (SDM) is a challenging remote-sensing task that involves establishing the relationship between geographical, environmental, and biophysical conditions and species presence. The goal of SDM is to produce large-scale range maps for species. Such maps can then be used to develop tools for modeling habitat suitability, predicting future distributions, etc. These tools are essential for biodiversity conservation and understanding the environmental effects of climate change [4].
Traditionally, species ranges were mapped using geographical and environmental features. The approaches were either limited to single-species models or modeled a fixed set of species. While the current approaches have their merits, we contend they fall short in accurately representing the inter-species relationships inherent in the taxonomic hierarchy. For example, species belonging to the same genera tend to be found in similar locations. They often share many common characteristics, including their general habitat preferences [45]. This pattern diminishes as one moves up in the taxonomic hierarchy. To capture these complex hierarchical relationships between the species, we use large language models (LLMs) to encode their taxonomic information in text (as shown in Figure 1) in the form of species-specific embeddings. Moreover, by leveraging the vast knowledge of LLMs, our approach enhances the identification of rare taxa [9].
Figure 1: Species Distribution Modeling using LLM Guidance. We use text prompts in the form of key-value pairs for training our species distribution model. This kind of prompt captures the taxonomical hierarchy of species. We use only the exhaustive prompt (d) for training while the rest are used for zero-shot predictions.
Previous approaches to SDM [30, 14, 26] tried to answer the question, “Which species are likely to be found at a given location?” However, to incorporate species-specific rich information into SDM, it is necessary to reformulate the SDM problem. Therefore, in this paper, we reformulate the SDM problem as “Which locations are likely for a given species to be found?” Assuming uniform prior over the species distribution [30, 26], this is an equivalent formulation. But by doing so, we can now incorporate a variety of information, including species-specific embeddings or any other metadata, which opens up new opportunities for more accurate modeling. Alternatively, this formulation makes it easier to incorporate unseen species that were not present in the training data.
@@
One of the principal challenges in training SDM models with presence-only data is properly evaluating the resulting species distributions. One approach is to compare to hand-crafted range maps, but such maps are difficult to create and not widely available with sufficient quality to be suitable. To this end, we propose a proximity-adjusted evaluation metric that penalizes predictions based on their distance to the closest true locations. We construct a One-Way Probability-Weighted Chamfer Distance (PWCD) metric, that measures the dissimilarity between the true and predicted species range maps, which is robust to small shifts. This enables an effective evaluation of SDMs using pixel-level range maps and crowdsourced data (where absence is not confirmed).
The key contribution of our work are as follows:
@@ 2 Related Works
Species distribution modeling (SDM) methods use environmental, geographic, and species observation data to predict species distribution across geographic space and time [17, 4]. In this context, we restrict ourselves to species range mapping over geographic space only. In the past, SDM relied on hand-crafted geographic and environmental features as input to capture the implicit spatial correlations in data [18, 32, 47]. These approaches were limited to one species per model due to their reliance on traditional supervised learning algorithms. With advancements in deep neural network architectures, end-to-end frameworks can model joint species distributions and complex implicit relations in data [6, 10, 14, 22, 27].
Another important aspect of SDM is the type of species occurrence data. Species occurrence data may exist in one of the two forms: presence-only and presence-absence. Collecting presence-absence observations is challenging, requiring expert surveys especially to confirm the absence of a given species in some region [31]. However, once collected, this kind of data can be easily fed into supervised learning frameworks, without requiring sophisticated modifications [20, 11, 21, 6, 44]. With no expert knowledge, presence-only data can be recorded when any species of interest is encountered. This way of collecting data does not require the data collector to confirm absences. However, since there are no negative samples, it is difficult to use standard supervised learning frameworks [10, 24, 8].
Most recent works formulate the problem of SDM using presence-only data as a multi-label classification task [30, 14]. Here, the goal of SDM is to predict the likelihood of species for a given location and environmental features. Many multi-label learning losses have been proposed in the literature that can easily be plugged in for this task. For instance, the AN-full (full assume negative) loss [30] randomly samples locations that are assumed to be negative during training. The ME-full (full maximum entropy) [54] tries to maximize the entropy of predictions of negative samples and at randomly sampled locations during training. Other losses such as ASL (asymmetric loss) [35], Hill loss [53], and RAL (robust asymmetric loss) [34] weigh the negative samples in a way that easy negative and potential false negative samples are down-weighted. Cole et al[14] evaluated the performance of SDMs trained with losses involving pseudo-negative sampling. We extend their work by evaluating the impact of training their model with the ASL and RAL losses.
@@
Finally, another line of work uses high-resolution satellite imagery at each species occurrence location to guide SDM methods [7, 13, 28, 29, 38, 43, 16, 39, 52]. For example, [43, 16] uses satellite imagery and geolocation to directly predict the likelihood over species while [38, 39] uses a contrastive learning framework to establish similarity between ground-level images of species and corresponding satellite imagery. These methods are limited in geographic scale due to the high-resolution nature of raw data [14]. Recent methods such as Range [15] proposed a retrieval augmented approach to learn geographic location representations using retrieved satellite imagery for a query location. Except [43, 16], all the works view this problem as an information retrieval task rather than a species prediction task.
@@ 3.1 Problem Definition
We consider presence-only data in the form of occurrence records: $\mathcal{D}=\{(l_{i},y_{i})|\>i=1,...,N\}$, where $l_{i}\in\mathbb{R}^{2}$ is the geolocation of each observation and $y_{i}\in[1,...,S]$ is the species of the observed organism. Additionally, a set of environmental features $e_{i}\in\mathbb{R}^{n}$ is included for each location $l_{i}$. Many previous studies [30, 26, 14] have tackled the task of SDM by training a neural network to model the conditional probability of p(y|l,e)p(y|\>l,e). This translates to the problem of “How likely a species can be found at a given location?” However, our approach involves modeling the probability of $p(r|\>y,l_{r},e_{r})$, which answers the question “Is it likely for a given species to be found at a given location?” Here, $r\in[1,...,L]$ is a set of discrete locations that are of interest when predicting the range of species belonging to some class. In this framework, we define the set of locations $r$ as a contiguous grid that is laid over the surface of the globe. The variables $l_{r}$ and $e_{r}$ represent fixed parameters in the model, namely the location vectors and environmental features defined at each location in $r$ respectively. It is easy to see that the input $y$ and the parameters $(l_{r},e_{r})$ can have any arbitrary representations. So, this new formulation allows for flexible encoding of both the species class and geographic location, compared to the previous formulation which only allowed flexibility in representing the geographic location.
@@ 3.2 Architecture Overview
Our objective is to train a neural network of the form: $f_{\theta}(y;l_{r},e_{r}):\mathbb{R}^{d}\rightarrow[0,1]^{L}$, where $d$ is the dimensionality of the species feature vector. $f_{\theta}$ is composed of three modules: 1) a location encoder network $f_{\alpha}$, 2) a species encoder network $g_{\beta}$, and 3) a multi-label classifier $h_{\gamma}$. See Figure 2 for a high-level architecture diagram. The location encoder network $f_{\alpha}$ consists of convolutional feature extractors and the Spherical Fourier Neural Operator (SFNO) [5] blocks that take as input the fixed parameters: $l_{r}$ and $e_{r}$. The network processes the global environmental covariates ($e_{r}$) to extract feature maps that represent the geography of a region. To incorporate geographic location-based features, we use $\sin$-$\cos$ location encodings ($l_{r}$) as learnable positional encodings for the SFNO blocks. The species encoder $g_{\beta}$ is a large language model that takes as input the text descriptions corresponding to a particular species of class $y$. This allows for a flexible representation of species-specific features in the form of natural language. The multi-label classifier $h_{\gamma}$ consists of a cross-attention network that fuses features from the SFNO and text encoder and performs multi-label classification using a linear layer. We call our model Language-Driven Species Distribution Model (LD-SDM).
@@ 3.3 Species-Specific Embeddings
One of the novel contributions of this paper is incorporating natural language descriptions of species for SDM. Many popular crowdsourcing science platforms, such as iNaturalist [48], contain not just occurrence records, but also valuable species-specific information. This information may include the complete taxonomic hierarchy of each species (such as class, order, family, etc.) or the IUCN red list category (such as endangered, near threatened, etc.). While often ignored, this fine-grained information could be incorporated to enhance the performance and interpretability of SDMs. In this work, we incorporate the taxonomic hierarchy of species into our SDM in the form of text. Each species is uniquely identified by its taxonomic classification ranks, which can be represented as key-value pairs. For instance, the key-value pairs that represent the wild duck are defined as: t = {class: Aves, order: Anseriformes, family: Anatidae, genus:  Anas, species: Anas platyrhynchos}.
Recently, Large Language Models (LLMs) have demonstrated excellent capabilities in representing implicit knowledge, linguistic structure, and hierarchical concepts when trained on massive amounts of data [1]. With no finetuning, pretrained LLMs have shown impressive zero-shot performance with zero in-domain data. In this work, we use the recently released, LLaMA-2 [46] to encode the species-specific key-value pairs as text. We experiment with three variants of LLaMA-2 of which the 70B model produced the best performance.
@@ 3.4 Loss Functions
Species distribution modeling is a special case of multi-label learning. Numerous losses for multi-label learning have been evaluated for SDM. Depending on the formulation of the SDM task, it may fall in the domain of Single Positive Multi-Label Learning (SPML), where only a single presence of species is known for a specific geographic location. We describe the loss functions considered in this work below and provide extensive experimentation to assess their impact in Section 4.4.
@@
$\mathcal{L}_{\text{AN-full}}(\hat{y},y)=-\frac{1}{S}\sum_{s=1}^{S}[\mathbbm{1}_{[y_{s}=1]}\lambda log(\hat{y}_{s})$ (1)
@@
$\mathcal{L}_{\text{ME-full}}(\hat{y},y)=-\frac{1}{S}\sum_{s=1}^{S}[\mathbbm{1}_{[y_{s}=1]}\lambda log(\hat{y}_{s})$ (2)
@@
where $\psi(\hat{y}_{s})=\delta-\hat{y}_{s}$ is the Hill loss term. $M$, NN, $\alpha_{n}$ and $\beta_{n}$ are balance parameters for positive and negative samples.
@@ 4.1 Implementation Details
As described in Section 3.2, we use SFNO [5] to extract geographic and environmental-specific features. We use two spectral and encoder layers in the SFNO. The encoder dimension used is 128 while the number of SFNO blocks are two. For the coordinate-only models, we drop the environmental features during the training. For the text encoder, we use frozen LLaMA-2 [46] (70B variant) and extract its encoder outputs for each species. Before training, we precompute and save the embeddings for each species in the disk. For the multi-label classifier, we use a single-layer multi-headed cross-attention network followed by a fully-connected layer with sigmoid activations. We compare LD-SDM with two state-of-art-models: SINR [14] and SIREN(SH) [37]. SINR learns an end-to-end neural representation of geographic locations that is used to predict the likelihood over species. SIREN(SH) uses a spherical harmonic representation of geographic locations which is then used to predict the likelihood over species. The rest of the details about our models and the implementation of state-of-the-art models are described in Appendix C.
@@ 4.2 Proximity-Aware Evaluation
When predicting species range, the evaluation process typically involves using range maps defined by experts or curated carefully [32, 47]. However, these maps can often be sparse and difficult to obtain. Alternatively, using crowdsourced species observation data is challenging since the absence of a given species is not recorded. Intuitively, models with false predictions close to their true locations are better than those with distant false predictions. This is because models in the former case learn spatially localized species distributions.
Figure 3: Performance of our spatially-aware metric against the MAP score. The figure depicts the inability of the MAP score to distinguish between range maps with varying spatial distributions of false positives. (a) Ground truth range map. (b) Predicted range map with false positives spatially distant from true positives. (c) Predicted range map with false positives spatially close to true positives. (d) Predicted range map without false positives.
Previous studies have used the area under the curve (AUC) [32, 47] and the mean average precision (MAP) [14] to evaluate SDMs. However, these metrics ignore the spatial distribution of predictions and weigh them equally, making them highly sensitive to spatial shifts in the predictions. This makes evaluating SDMs a challenging task since it requires the comparison of spatially distributed probability values. Pixel-wise comparison methods are not well-suited for this task since they are susceptible to spatial prediction shifts.
@@
A well-known spatially explicit distance used for comparing any two distributions is the Earth-Mover (Wasserstein) Distance. Due to its computational requirements and infeasibility on large-scale data points, Chamfer Distance is a popularly used proxy [3]. However, it has only been developed for point clouds and there doesn’t exist a chamfer distance measure to compare two spatial probability distributions. Inspired by [50], we modify their Density-Aware Chamfer Distance (DCD) metric that works for comparing a predicted probability distribution map ($P_{1}$) with a binary probability map ($P_{2}$) representing the ground truth. We add a likelihood probability term which weights the distance term. We call this metric One-Way Probability Weighted Chamfer Distance (PWCD), which measures the false positive rate in a proximity-aware manner. This metric is between [0,1] and could also be used to compute the number of false positives instead of the rate. It is given by:
[…]
4.3 Training and Evaluation
We focus on bird species occurrence data which is abundantly available yet challenging to model as compared to other species due to complex migration patterns. In Table 1, we describe the overview of each dataset compiled in this work. We compiled 131.7 million observations of 4141 species of birds using the GBIF (gbif.org) platform which includes data from various community science platforms. This data is collected for the years 2012-2020. For evaluation, we use curated research-grade observations from the iNaturalist [48] and eBird [41] for 2021 and 2022. These contain 17.9 million and 17.2 million observations respectively. Using the same platform, we created a dataset of 84 species that are unseen (not present in training data) and rare (<< 1000 observations) for zero-shot performance evaluation. This contains a total of 26k observations. These platforms are useful since misclassification errors are highly mitigated due to community consensus [26], hence being a strong representative of the ground truth. Further, we use the same procedure as described in [14] for the geo-feature regression task.
Table 1: Number of observations in each dataset compiled in this work.
Dataset Year #Observations #species
Train 2012-2020 131.7M 4141
Test 2021 17.9M 3960
Test 2022 17.2M 4010
Zero-Shot 2021-2022 26k 84